# Magneto: Combining Small and Large Language Models for Schema Matching

### v1

@@ Abstract.
Recent advances in language models opened new opportunities to address complex schema matching tasks. Schema matching approaches have been proposed that demonstrate the usefulness of language models, but they have also uncovered important limitations: Small language models (SLMs) require training data (which can be both expensive and challenging to obtain), and large language models (LLMs) often incur high computational costs and must deal with constraints imposed by context windows. We present Magneto, a cost-effective and accurate solution for schema matching that combines the advantages of SLMs and LLMs to address their limitations. By structuring the schema matching pipeline in two phases, retrieval and reranking, Magneto can use computationally efficient SLM-based strategies to derive candidate matches which can then be reranked by LLMs, thus making it possible to reduce runtime without compromising matching accuracy. We propose a self-supervised approach to fine-tune SLMs which uses LLMs to generate syntactically diverse training data, and prompting strategies that are effective for reranking. We also introduce a new benchmark, developed in collaboration with domain experts, which includes real biomedical datasets and presents new challenges to schema matching methods. Through a detailed experimental evaluation, using both our new and existing benchmarks, we show that Magneto is scalable and attains high accuracy for datasets from different domains.
@@
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
@@
The source code, data, and/or other artifacts have been made available at https://github.com/VIDA-NYU/data-harmonization-benchmark and https://github.com/VIDA-NYU/data-integration-eval.
@@ 1. Introduction
Proteogenomic analysis of endometrial cancer tumors: In the context of the National Cancer Institute’s Clinical Proteomic Tumor Analysis Consortium (CPTAC) (cptac, 2024), Li et al. (2023) carried out a comprehensive proteogenomic analysis of endometrial cancer tumors. They pooled data from ten studies (available as supplementary material in the original primary research papers) that cover multiple patient cohorts and cancer types. To facilitate the integration and analysis, they mapped each dataset into the GDC standard, a data standard set by the National Cancer Institute’s Genomic Data Commons (GDC) for cancer genomic data (Heath et al., 2021; National Cancer Institute, 2024a). Even though the studies had been carried out by members of the CPTAC effort, the datasets containing patient case and sample data used many different representations for variable names and values. Integrating these data required a substantial effort to match variables from each dataset source schemata to the target GDC format, which encompasses over 700 attributes.
@@
Even though the schema matching (SM) problem has been extensively studied in the data management literature (Doan et al., 2012; Cafarella et al., 2009; Miller, 2018; Koutras et al., 2021), we lack solutions that are general and work well for data in the wild. Schema matching still requires a time-consuming, manual curation process for complex matching tasks like the one described above (Adhikari et al., 2021; Cheng et al., 2024). Matching is particularly challenging in the presence of ambiguity and heterogeneity in the representation of attributes and values (see Table 1). SM approaches that rely on attribute names, data types and values for similarity assessments are likely to fail for such matches. As a point of reference, we assessed the effectiveness of state-of-the-art schema matching strategies for the biomedical datasets from Example 1.1. As shown in Figure 1, they perform poorly: the best technique achieves at most 0.45 of mean reciprocal rank (MRR) and requires powerful computational resources to run. We discuss these results in detail in Section 6.
Figure 1. Trade-off between runtime and accuracy measured as the mean reciprocal ranking (MRR) for different schema matchers for the matching task of Example 1.1. Magneto variants (in orange) achieve higher MRR and better accuracy-runtime trade-offs than traditional methods (in blue).
@@
Determining columns correspondences may require knowledge beyond the schema and contents of a table. Table 1 shows some possible matches for the proteogenomic analysis. Sometimes, selecting the correct match is difficult even for subject matter experts. For example, for the attribute patient_age from one of the datasets, there are at least three plausible matches in GDC: age_at_diagnosis, days_to_birth, and age_at_index. Without additional context, it is difficult to determine the correct match.
@@
Schema Matching and Language Models. There has been renewed interest in data integration given the new capabilities introduced by language models (Feuer et al., 2024; Narayan et al., 2022). For schema matching, approaches have been proposed that have shown promising results (Castro Fernandez et al., 2018; Cappuzzo et al., 2020; Du et al., 2024; Tu et al., 2023). A key challenge in schema matching is the estimation similarities between two columns. Pre-trained Language Models (PLMs) ,referred to as SLMs in this paper to distinguish them from large language models (LLMs), support the creation of representations, or embeddings, for columns enriched with semantic information. The similarity between two embeddings can serve as a proxy for column-matching scores (Castro Fernandez et al., 2018; Cappuzzo et al., 2020). SLMs have also been fine-tuned for schema matching (Du et al., 2024) and general matching tasks (Tu et al., 2023), leading to accuracy gains compared to the general models.
LLMs are trained using large, generic data corpora and thus contain knowledge that can assist in obtaining additional semantics necessary to identify matches. Prompting strategies combined with fine-tuned models have been shown effective to improve table understanding and help in integration tasks (Li et al., 2024a).
While prior work has shown the usefulness of SLMs and LLMs for SM, they also present significant practical practical challenges:
  • $\bullet$
  • $\bullet$
    Challenge 2: LLMs usually do not need fine-tuning, but they have a fixed context window: if a prompt exceeds this limit, it may need to be truncated before it is sent to the model. Since small context windows are common in LLMs, this can lead to losing critical context and limit their use for large schemata and tables. While recent LLMs are starting to support larger windows, their cost scales with input and output size. Besides, recent studies have shown that the accuracy of LLMs can significantly decrease when given long prompts for complex tasks (Li et al., 2024b). Last but not least, calls to LLM APIs can have high latency, especially for large prompts.
Table 1. Column names and values for candidate matches between columns in source tables and the GDC format. These illustrate challenges involved in identifying matches in the presence of heterogeneity in the schemata. Note: GDC does not specify values for numerical attributes such as age.
@@
Path_Stage_Reg_Lymph_Nodes-pN: pN1 (FIGO IIIC1), pN0, pNX ajcc_pathologic_n: N1, N0, NX
uicc_pathologic_n: N1, N0, NX
@@
Our Approach. We introduce Magneto, a framework that combines small (pre-trained) language models (SLMs) and large language models (LLMs) to derive cost-effective and general solutions for SM. As illustrated in Figure 2, Magneto is structured in two phases: candidate retrieval selects a subset of the possible matches; and reranker, which ranks the candidates to make it easier for users to examine and select matches. We use this framework to explore three new strategies that use SLMs and LLMs: Using SLM as the candidate retriever and LLMs as the reranker; an automated, LLM-powered fine-tuning process for SLMs; and a combination of these. We propose new methods that address the context window and cost challenges associated with LLMs and the creation of training data to fine-tune SLMs.
To address Challenge 1, we leverage LLMs to automate the process of fine-tuning SLMs. Instead of relying on manually created training data and structure-based augmentation (e.g., row shuffling and sampling (Fan et al., 2023; Du et al., 2024)), we use LLMs to derive data. Using LLMs, we can add syntactic diversity and capture different representations for column names and their values.
@@
We use a multi-pronged approach to address Challenge 2. By structuring the matching process in two phases, Magneto can use cheaper SLM-based methods for filtering candidates, thus reducing the number of matches that need to be checked by more costly rerankers, e.g., LLM-based rerankers, thereby reducing the cost and latency. We must also ensure that all necessary details are included in the prompt while staying within the context window limit. Thus we need to create representative samples the contents of the columns in candidate matches to be ranked by the LLM. In addition, LLMs are designed for textual data. To use them for tabular data, we must serialize the column content. Selecting the right serialization strategy is still an open research problem that has attracted substantial attention (Fang et al., 2024; Dong and Wang, 2024). We explore different alternatives for both sampling and serialization.
@@
Our ultimate goal is to devise strategies that have high accuracy and low cost. However, there is often a trade-off: high-accuracy is associated with higher costs and running times. Figure 1 illustrates the trade-offs between runtime and accuracy for existing schema matching algorithms and Magneto. Traditional methods tend to cluster at relatively high accuracy with long runtimes or relatively low runtimes with reduced accuracy. In contrast, Magneto variants strike a balance across both dimensions and have the advantage of requiring no prior training data.
@@
We conduct an extensive experimental study of Magneto on various datasets from two benchmarks. The first is a new benchmark, that we developed in collaboration with domain experts, which includes real biomedical datasets. Our experiments show that many existing SM solutions struggle with this benchmark, particularly given the complexity and syntactic variability present in its datasets. In contrast, Magneto demonstrates superior accuracy and generally faster runtimes. We also evaluate Magneto on the Valentine benchmark (Koutras et al., 2021) and observe that it also performs well for the Valentine datasets which cover a variety of domains.
Figure 2. Magneto overview. Given source and target tables, the framework identifies matches in two phases: A small language model, automatically fine-tuned using LLM-generated data, retrieves and ranks candidate matches, which are then assessed and reranked by an LLM. This process makes it possible to efficiently identify matches and generalize across domains and datasets. Different schema matching strategies can be derived by adopting different combinations of methods for these phases.
@@
    We introduce the Magneto framework that effectively combines SLMs and LLMs and allows the creation of SM strategies that attain a balance between accuracy and runtime for diverse data and SM tasks (Section 2).
@@
    We propose an LLM-powered method that automatically generates training data needed to fine-tune SLMs, making it possible to effectively use these low-cost models for specific SM tasks. We also present a fine-tuning pipeline that uses contrastive learning and, by combining triplet loss and online triplet mining, effectively utilizes the generated data to enhance the discriminative power of column embeddings for SM (Section 3).
@@
    Unlike previous approaches that use LLMs for SM, we show that reranking candidate matches with LLMs, derived from less costly methods, achieves high accuracy at lower costs. (Section 4).
@@
    In collaboration with biomedical researchers, we created a new benchmark that represents a real schema matching effort (Li et al., 2023), encompassing characteristics not present in existing benchmarks and presents new challenges for SM methods (Section 5).
@@
    We perform an extensive experimental evaluation of different strategies derived using our framework, comparing them against SM methods over datasets that cover a wide range of domains. We also carry out ablation studies to assess the effectiveness of different choices for column sampling and serialization. (Section 6).
@@ 2. Motivation and Solution Overview
Biomedical Data Integration: Opportunities and Challenges. There have been substantial investments in infrastructure to share biomedical data (National Institutes of Health, 2020). Today, many datasets are available in repositories and published with scientific articles (pdc, 2023; National Cancer Institute, 2024a). The ability to pool these data can advance research in multiple diseases and across different populations. However, established technologies are limited in their ability to integrate these data. Our research was motivated by this challenge as part of a collaboration with biomedical researchers in the context of the ARPA-H Biomedical Data Fabric (BDF) program (ARPA-H, 2024), which aims to develop data integration and usability technologies.
We conducted formative interviews with biomedical experts and found that they rely heavily on manual processes and dataset-specific scripts to perform schema matching (Mertins et al., 2016; Li et al., 2023; Cheng et al., 2024). The example described in Section 1 illustrates some of the difficulties encountered in one study. Besides the time required, the process they follow is error-prone and difficult to reproduce. As data integration is a key step in most of their studies, they would benefit from automated (and semi-automated) techniques that streamline integration and are also transparent and reproducible.
We have identified two important requirements based on the researchers’ needs. They need approaches that: generalize across different kinds of data (e.g., genomics, proteomics, electronic health records), for different diseases (e.g., different types of cancers, auto-immune and rare diseases), from many sources (e.g., publications, hospitals, data commons); and that facilitate curation, since in many cases it’s difficult (even for experts) to determine correct matches. These requirements have guided the design of Magneto.
@@ 2.1. Definitions and Evaluation Metrics
Before we describe our approach, we introduce the terminology used in the remainder of this paper. We formally define the problem we are addressing and the performance metrics we use to evaluate the methods we propose for schema matching.
@@
Matching algorithms often associate a score with each match they derive. These scores are used to generate ranked lists containing the derived matches, which help users explore matches by prioritizing the highest-scoring candidates. These lists provide a global ranking of best matches $(A_{i},B_{j})$ among all pairs of possible matches of a given pair of tables S and T or a per-column ranking of the best matches $B_{j}$ for a given source attribute $A_{i}\in\textbf{S}$. Thus, a common approach to evaluate schema matching methods is to assess their ability to produce high-quality ranked lists of matches (Koutras et al., 2021; Du et al., 2024). We use two measures in our evaluation, Mean Reciprocal Rank (MRR) and Recall at Ground-Truth Size (Recall@GT), which capture different aspects of matching methods: user experience and result completeness.
@@
(Mean Reciprocal Rank) Let $matches[A]$ denote the ranked list of matches produced by a schema-matching algorithm for the source column $A\in\textbf{S}$, and $r_{A}$ be the position of the first correct target column $B$ within the ordered list $matches[A]$. The reciprocal rank (RR) of an individual column $A$ is the multiplicative inverse of the rank, i.e., $\frac{1}{r_{A}}$. For instance, the RR is $1$ if the match is found in the first place, $\frac{1}{2}$ for the second place, $\frac{1}{3}$ for the third place, and so on. The mean reciprocal rank (MRR) for a table S is the average RR over the subset of columns $\textbf{S}^{\prime}$ that contain a correct match in the ground truth:
@@
Intuitively, MMR measures how long it takes to find the first relevant match when examining the ranked list of matches. ∎
MRR is commonly used to evaluate methods that produce a ranked list of items, including the evaluation of search engine results and responses derived by question-answering systems (Schütze et al., 2008; Voorhees et al., 1999). In short, MRR evaluates the quality of the ranking. The prioritization of the top-ranked correct result aligns with an enhanced user experience. In our problem, this corresponds to ranking the most relevant match for a given source column as highly as possible: a high MRR score indicates that a user can find the correct match more easily when judging multiple matches for a column.
We also use the recall at ground truth size (Recall@GT), a standard measure used in recent schema matching works (Koutras et al., 2021). In contrast to MRR, Recall@GT is computed over a global ranking of the matches, i.e., instead of considering the ranking for each source column, it considers a merged list of all matches produced.

### v2

@@ Abstract.
Recent advances in language models (LMs) open new opportunities for schema matching (SM). Recent approaches have shown their potential and key limitations: while small LMs (SLMs) require costly, difficult-to-obtain training data, large LMs (LLMs) demand significant computational resources and face context window constraints. We present Magneto, a cost-effective and accurate solution for SM that combines the advantages of SLMs and LLMs to address their limitations. By structuring the SM pipeline in two phases, retrieval and reranking, Magneto can use computationally efficient SLM-based strategies to derive candidate matches which can then be reranked by LLMs, thus making it possible to reduce runtime while improving matching accuracy. We propose (1) a self-supervised approach to fine-tune SLMs which uses LLMs to generate syntactically diverse training data, and (2) prompting strategies that are effective for reranking. We also introduce a new benchmark, developed in collaboration with domain experts, which includes real biomedical datasets and presents new challenges for SM methods. Through a detailed experimental evaluation, using both our new and existing benchmarks, we show that Magneto is scalable and attains high accuracy for datasets from different domains.
@@
PVLDB, 18(8): 2681 - 2694, 2025.
doi:10.14778/3742728.3742757 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 8 ISSN 2150-8097.
doi:10.14778/3742728.3742757
@@
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/VIDA-NYU/magneto-matcher.
@@ 1. Introduction
In the context of the National Cancer Institute’s Clinical Proteomic Tumor Analysis Consortium (CPTAC) (cptac, 2024), Li et al. (2023) carried out a comprehensive proteogenomic analysis of cancer tumors. They collected data from ten studies (published as supplementary material in research papers) that cover multiple patient cohorts and cancer types. To facilitate their analysis, they mapped each dataset into the GDC standard, a data model set by the National Cancer Institute’s Genomic Data Commons (GDC) for cancer genomic data (Heath et al., 2021; National Cancer Institute, 2024a). Even though the studies had been carried out by members of the CPTAC effort, the datasets containing patient case and sample data used different representations for variable names and values. Integrating these data required a substantial effort to match variables from each dataset source schemata to the target GDC format, which encompasses over 700 attributes. ∎
@@
Even though the schema matching problem has been extensively studied (Doan et al., 2012; Cafarella et al., 2009; Miller, 2018; Koutras et al., 2021), matching still requires a time-consuming, manual curation process for complex tasks, which like the one described above, involves ambiguity and heterogeneity in the representation of attributes and values (see Table 1). Schema matching approaches that rely on attribute names, data types and values for similarity assessments are likely to fail for such matches. As a point of reference, we assessed the effectiveness of state-of-the-art strategies for the biomedical datasets from Example 1.1. As shown in Figure 1, they perform poorly: the best technique achieves at most 0.45 of mean reciprocal rank (MRR) and incurs high computational costs. We discuss these results in detail in Section 6.
@@
Determining correspondences between columns may require knowledge beyond the schema and contents of a table. Table 1 shows some possible matches for the proteogenomic analysis. Sometimes, selecting the correct match is difficult even for subject matter experts. For example, for the attribute patient_age from one of the datasets, there are at least three plausible matches in GDC: age_at_diagnosis, days_to_birth, and age_at_index. Without additional context, it is difficult to determine the correct match.
@@
Schema Matching and Language Models. Renewed interest in data integration has emerged due to the capabilities of language models (Feuer et al., 2024; Narayan et al., 2022). For schema matching, promising approaches have been proposed (Castro Fernandez et al., 2018; Cappuzzo et al., 2020; Du et al., 2024; Tu et al., 2023). A key challenge in schema matching is estimating the similarity between two columns. Pre-trained Language Models (PLMs), referred to as SLMs in this paper to distinguish them from large language models (LLMs), create column representations (or embeddings) enriched with semantic information. The similarity between two embeddings serve as a proxy for column-matching scores (Castro Fernandez et al., 2018; Cappuzzo et al., 2020). SLMs have also been fine-tuned for schema matching (Du et al., 2024) and general matching tasks (Tu et al., 2023).
Table 1. Examples of candidate matches between source tables and the GDC that highlight schema heterogeneity.
@@
LLMs, in contrast, are trained using large, generic data corpora and thus contain knowledge that can assist in obtaining additional semantics necessary to identify matches. Prompting strategies combined with fine-tuned models have been shown effective to improve table understanding and help in integration tasks (Li et al., 2024a).
While prior work has shown the usefulness of SLMs and LLMs for schema matching, they also present significant practical challenges:
  • $\bullet$
  • $\bullet$
    Challenge 2: LLMs usually do not need fine-tuning, but face constraints due to fixed context windows, requiring truncation of large prompts and potential loss crucial information. While some models offer larger windows, cost scales with input and output size. Furthermore, accuracy can decline with long prompts and API calls incur high latency, especially for large inputs (Li et al., 2024b).
@@
Our Approach. We introduce Magneto, a framework that combines SLMs and LLMs to derive cost-effective and general schema matching solutions. As illustrated in Figure 2, Magneto is structured in two phases: candidate retrieval selects a subset of the possible matches; and reranker, which ranks the candidates to make it easier for users to examine and select matches. To address Challenge 1, Magneto leverages LLMs to automate SLMs fine-tuning. Instead of relying on manually created training data and structure-based augmentation (e.g., row shuffling and sampling (Fan et al., 2023; Du et al., 2024)), we use LLMs to derive data. Using LLMs, we can add syntactic diversity and capture different representations for column names and their values. Magneto addresses Challenge 2 by using cheaper SLM-based methods for finding candidates thereby reducing the number of matches that need to be checked by more costly LLM-based rerankers. To ensure that all necessary details are included in the prompt while staying within the context window limit, we need to create representative samples of the columns in candidate matches to be ranked by the LLM. In addition, since LLMs are designed for textual data, we must serialize the column content. Selecting the right serialization strategy is still an open research problem that has attracted substantial attention (Fang et al., 2024; Dong and Wang, 2024). We explore different alternatives for both sampling and serialization.
@@
Our goal with Magneto is to achieve high accuracy for schema matching at a low cost. Figure 1 illustrates the trade-offs between runtime and accuracy for existing schema matching algorithms and different variants of Magneto. Note that traditional methods tend to cluster at relatively high accuracy with long runtimes or relatively low runtimes with reduced accuracy. In contrast, Magneto variants strike a balance across both dimensions and have the advantage of requiring no training data.
@@
We conduct an extensive experimental study of Magneto on various datasets from two benchmarks. The first is a new benchmark, that we developed in collaboration with domain experts, which includes real biomedical datasets. Our experiments show that many existing schema matching solutions struggle with this benchmark, particularly given the complexity and syntactic variability present in its datasets. In contrast, Magneto demonstrates superior accuracy and generally faster runtimes. We also evaluate Magneto on the Valentine benchmark (Koutras et al., 2021) and observe that it also performs well for the Valentine datasets which cover a variety of domains.
Figure 1. Trade-off between runtime and accuracy (using MRR) for different schema matchers in the task of Example 1.1. Magneto variants (orange) outperform traditional methods (blue) in MRR and accuracy-runtime trade-off.
@@
    We introduce the Magneto framework that effectively combines SLMs and LLMs and allows the creation of schema matching strategies that attain a balance between accuracy and runtime for diverse data and tasks (Section 2).
@@
    We propose an LLM-powered method to generate training data for fine-tuning SLMs for schema matching tasks and a contrastive learning pipeline using triplet loss and online triplet mining to enhance column embedding distinction (Section 3).
@@
    Unlike previous approaches that use LLMs for schema matching, we show that reranking candidate matches with LLMs, derived from less costly methods, achieves high accuracy at a lower cost. (Section 4).
@@
    In collaboration with biomedical researchers, we created a new benchmark that represents a real schema matching effort (Li et al., 2023), encompassing characteristics not present in existing benchmarks and introducing new challenges for SM methods (Section 5).
@@
    We perform an extensive experimental evaluation of different strategies derived using our framework, comparing them against SM methods over datasets that cover a wide range of domains. We also carry out ablation studies to assess the effectiveness of different choices for column sampling and serialization (Section 6).
Figure 2. Magneto takes source and target tables and identifies matches in two phases: an SLM retrieves and ranks candidates, then an LLM assesses and reranks them. This enables efficient matching and cross-domain generalization, with customizable strategies based on method combinations.
@@ 2. Motivation and Solution Overview
Biomedical Data Integration: Opportunities and Challenges. Due to substantial investments made in infrastructure to share biomedical data (National Institutes of Health, 2020), many datasets are now stored in repositories with related research articles (pdc, 2023; National Cancer Institute, 2024a). Pooling these data can advance research across various diseases and populations, but current technologies are limited in their ability to integrate these data. Our research was motivated by this challenge as part of a collaboration with biomedical researchers in the context of the ARPA-H Biomedical Data Fabric (BDF) program (ARPA-H, 2024), which aims to develop usable data integration methods and tools.
Our interviews with biomedical experts revealed that current data integration practices rely heavily on manual processes and dataset-specific scripts for schema matching (Mertins et al., 2016; Li et al., 2023; Cheng et al., 2024). As illustrated in Example 1.1, this approach is both error-prone and difficult to reproduce, creating significant barriers to biomedical research.
Based on these findings, we identified two critical requirements that guided the design of Magneto. First, researchers need approaches that generalize across the diverse landscape of biomedical data—spanning different data types (genomics, proteomics, electronic health records), disease categories (various cancers, autoimmune and rare diseases), and data sources (publications, hospitals, data commons). Second, given the complexity of biomedical schemas, these approaches must facilitate curation, as even subject matter experts often struggle to determine correct matches.
@@ 2.1. Definitions and Evaluation Metrics
Before describing our approach, we introduce the notation, schema matching definition, and evaluation metrics used in this paper.
@@
Matching algorithms often associate a score with each match they derive. These scores are used to generate ranked lists containing the derived matches, which help users explore matches by prioritizing the highest-scoring candidates. These lists provide a global ranking of best matches $(A_{i},B_{j})$ among all pairs of possible matches of a given pair of tables S and T or a per-column ranking of the best matches $B_{j}$ for a given source attribute $A_{i}\in\textbf{S}$. Thus, a common approach to evaluate schema matching methods is to assess their ability to produce high-quality ranked lists of matches (Koutras et al., 2021; Du et al., 2024). We use two evaluation metrics: Mean Reciprocal Rank (MRR) and Recall at Ground-Truth Size (Recall@GT), as detailed next.
@@
(Mean Reciprocal Rank) Let $matches[A]$ denote the ranked list of matches produced by a schema-matching algorithm for the source column $A\in\textbf{S}$, and $r_{A}$ be the position of the first correct target column $B$ within the ordered list $matches[A]$. The reciprocal rank (RR) of an individual column $A$ is the multiplicative inverse of the rank, i.e., $\frac{1}{r_{A}}$. The mean reciprocal rank (MRR) for a table S is the average RR over the subset of columns $\textbf{S}^{\prime}$ that contain a correct match in the ground truth:
@@
Intuitively, MRR measures how long it takes to find the first relevant match when examining the ranked list of matches. ∎
MRR is a standard evaluation metric for ranked lists in search engines and question-answering systems (Schütze et al., 2008; Voorhees et al., 1999). For schema matching, high-quality results correspond to ranking the most relevant match for each source column as highly as possible. A high MRR score therefore indicates that users can more easily identify correct matches when evaluating candidate matches for a given column.
We also use the recall at ground truth size (Recall@GT), a standard measure used in recent schema matching literature (Koutras et al., 2021). Unlike MRR, which evaluates rankings per source column, Recall@GT operates on a global ranking that merges all candidate matches across columns into a single list.